Skip to main contentSkip to solution

Probabilities of solving a specific mathematical problem independently by three students are 12,13\frac{1}{2},\frac{1}{3} and 14\frac{1}{4}. If all the three students try to solve the problem independently, the probability that the problem is solved, is:

Solution

✅ Correct Option: 3

The problem stays unsolved only if all three fail: 12×23×34=14\frac12\times\frac23\times\frac34=\frac14. Hence P(solved) =1−14=34=1-\frac14=\frac34.

Keyboard Shortcuts

  • Left arrow: Previous question
  • Right arrow: Next question
  • S key: Jump to solution
  • Q key: Jump to question